the states C and D spend a total of $55.9 million for touism the state C spends 3.9 million more than state D find the amount that each states spends
Answer by Cromlix(4381)

Hi there,
states C and D spend a total of $55.9 million
the state C spends 3.9 million more than state D
State D = 'x'
State C = x + 3.9
C + D = $55.9 million
(x + 3.9) + x = $55.9
Collect like terms
2x + 3.9 = $55.9
2x = $55.9 - 3.9
2x = $52
x = $26
State C spends $26 + $3.9
= $29.9 million
State D spends $26 million
Hope this helps:-)